<<TableOfContents(3)>>

= Agenda/Minutes =
[[https://zoom.us/j/203743987|Zoom Meeting]]

----------
= CRY+VIS Interface Meeting =
Date: 2022/11/11 13:00-13:20

Participants: Takahashi, Mitsuhashi, Sato, Uchiyama, Ushiba, Ikeda

[[https://paper.dropbox.com/folder/show/AcceptanceTest-e.1gg8YzoPEhbTkrhvQwJ2zzo3dHVh5sJGieDgbm2n3h60mGqotBeE|Acceptance check document]]

== Status ==
 * Type-A: Sensor correction in EY, EX, IX, IY. Drift of the F0 in IY. Update of the MN Pitch control in EY.
 * Type-B: Script to calculate the sensor correction.
 * Type-Bp: 
 * Type-C: Trouble and recovery of the OSTM suspension. Health check in the next week.

== Other items ==
 * Replacement of cables: EXT remains.
 * Oplev: 
 * ACC: 
 * Commissioning: 
 * Vacuum system: GV opening in EYT after 11/15.
 * Heater control: 
 * TMS: 

== Discussion ==

----------
= VIS Meeting =
Date: 2022/11/11 13:30-13:40

Participants: Takahashi, Uchiyama, Sato, Tanaka, Mitsuhashi, Hirata, Lucia

== Summary Report ==
=== Type-A ===
 * The F0 in IY had a large drift, but it was recovered with room temperature control.
 * Test of the new IP blending with the Geophones and the LVDTs sensor correction in EY.
 * Check of the sensor correction performance in EX, IX, and IY.

=== Type-B ===
=== Type-Bp ===
=== Type-C ===
 * During the adjustment of the OSTM suspension, two magnets (H2, H3) on the mirror dropped off. We made the in-situ gluing jig for the magnets and glued them successfully.
 * Finally the OSTM was aligned not using the picomotor but rotating the suspension frame in yaw.

=== Electronics ===
=== Control ===

== Schedule ==
 * Type-A: 
 * Type-B:
 * Type-Bp: 
 * Type-C: Finalization of the OMMT1/2/OSTM suspensions.

== Safety ==
== Discussion ==
== Travel Plans ==
(Date in Kamioka)

The week from 11/14

 * Takahashi: 11/16-18
 * Hirata: 11/15-18
 * Sato: 11/15-18

The week from 11/21

 * Takahashi: 
 * Sato : 
 * Hirata: 

== Next meeting ==
 * 2022/11/18(FRI) 13:30